### v4

@@
Abstract   New liquid drop model with the iso-scalar volume and surface energy terms is applied to reproduce experimentally known masses of nuclei with number of protons and neutrons larger or equal to twenty. The ground-state microscopic energy corrections are taken into account. Even though the model contains only six adjustable parameters, the quality of mass reproduction is good and comparable with other contemporary mass models. Also, the fission barrier heights of actinide nuclei evaluated using the topographical theorem of Myers and Świa̧tecki are close to the data.
Keywords: mac-mic model, nuclear masses, fission barrier heights, scission point properties
@@ 1 Introduction
Mass is the most fundamental property of atomic nuclei for its relevance in understanding not only the nuclear structure and the complex interaction in nuclei, but also the path of the synthesis of heavy nuclei in astrophysics [1]. In the super heavy region, the mass data also essential in the exploration of synthesizing super-heavy nuclide in terrestrial laboratory. Experimentally, the cooling store ring and the ion trap both provide precise measurement of nuclear mass in modern age [2, 3, 4, 5, 6]. So far, approximately 3000-4000 nuclides are measured while more than 9000 nuclides are believed to exist according to theoretic model calculations [7] . Descriptions and predictions of nuclear mass are of high theoretic interests and of significance in the calculation of reaction network in stellar environment including r-process [8]. Many models have been developed for this purpose. Roughly speaking, the models are classified in three groups. i) The macro-micro models, including finite range droplet model (FRDM) [9] and Weizsäcker-Skyrme (WS) models [10, 11, 12], ii) The pure micro models including Hartree-Fock-Bogoliubov (HFB) models [13] and the covariant density-functional theory (CDFT) [14] etc. and iii) The local mass models which predict the atomic nucleus mass through the relationship of the masses of neighboring nuclides by taking into account different physical constraints [15, 16], such as Garvey-Kelson (GK) relations [17, 18] and the isobaric multiplet mass equation (IMME) [19] etc. Meanwhile, many efforts have been taken to evaluate and compare the predictive power of the models [20]. With the applications of the modern machine learning or neural network techniques, the predictive power of the models are further improved [21, 22, 23, 24, 25, 26]. Up to now, the residues of the nucleus mass reach the level of about 200 keV or better.
@@
A modern version of the LD model was proposed in 1966 by Myers and Świa̧tecki [32]. They have shown that the LD energy enriched by the shell and pairing effects can describe well the binding energies and quadrupole moments of known nuclei and gives a reasonable description of the fission-barrier height of heavy nuclei (see also [33]). Unfortunately, neither the Myers and Świa̧tecki LD formula nor its refined version called the droplet model could adequately reproduce the barrier heights of medium-heavy and lighter nuclei [34]. In addition, it was shown by von Groote and Hilf [35] that a further correction to the LD model, namely the curvature term, did not change much in this respect. Due to these results further development of the nuclear LD model was stopped in practice for more than three decades. Other much more complicated models like the droplet, Yukawa-folded (YF), Yukawa-plus-Exponential (YpE), or Finite-Range Droplet Models (FRDM) have been used to obtain within the so-called macroscopic-microscopic (mac-mic) approximation [32] of the binding energies and the fission barrier heights (for overview look, e.g., Ref. [36]). A more detailed historical review of the LD models can be found, e.g., in Ref. [37].
@@ 2 Theoretical model
$\begin{array}[]{ll}E_{\rm ISLD}(Z,A;{\rm sph})&=\displaystyle-a_{\rm vol}A\,\left(1-4\kappa_{\rm vol}\frac{T(T+1)}{A^{2}}\right)\\[12.91663pt] &\displaystyle+a_{\rm sur}A^{2/3}\,\left(1-4\kappa_{\rm sur}\frac{T(T+1)}{A^{2}}\right)\\[12.91663pt] &\displaystyle+{3\over 5}e^{2}\frac{Z(Z-1)}{A^{1/3}}+E_{\rm odd}(Z,A)\\[12.91663pt] &\displaystyle-{3\over 4}e^{2}\left({3\over 2\pi}\right)^{2/3}\frac{Z^{4/3}}{r_{0}A^{1/3}}~{}.\end{array}$ (2)
@@
$E_{\rm odd}(Z,A)=\left\{\begin{array}[]{ll}{\cal D}/Z^{1/3}{}&{\rm for~{}Z~{}odd~{},}\\ {\cal D}/N^{1/3}{}&{\rm for~{}N~{}odd~{},}\\ 0&{\rm for~{}Z~{}and~{}N~{}even~{}.}\end{array}\right.$ (3)
@@
Note that the linear in isospin term in the ISLD model (2) corresponds to the Wigner (or congruence) energy present in typical LD-like models (confer, e.g., [32, 39]). In addition, in the ISLD model, the deformation dependence of this linear in $|N-Z|$ is well defined, whereas, in Ref. [39], one has to add an additional phenomenological deformation dependence in order to obtain a doubling of the congruence at the scission point, when two fission fragment nuclei are born. The only free, i.e., adjustable, parameters of the ISLD model are: $a_{\rm vol},\,\kappa_{\rm vol},\,a_{\rm sur},\,\kappa_{\rm sur},\,r_{0}$, and ${\cal D}$.
@@
and it is equal to the difference between the ground-state energy of the nucleus and the spherical macroscopic energy.
Taking the above microscopic energy correction from the tables [54], we make an approximation. Namely, we assume that the macroscopic energy evaluated within the FRLDM at small deformations corresponding to the ground state of nuclei is close the that obtained using the ISLD model. Such approximation is acceptable when the macroscopic fission barriers obtained in both models are close to each other, which is the case here. A similar approximation was also made in Refs. [39, 38, 44].
The Fourier-over-Spheroid (FoS) shape parametrization [48, 49] has been shown to reproduce nuclear shapes very close to the optimal shapes obtained by the Strutinsky variational procedure [50] and allows to evaluate precisely the liquid drop energy at the saddle point. Knowing this energy, one can estimate the fission barrier height ($V_{\rm sadd}$) with the help of the Myers and Świa̧tecki topographical theorem [39]
@@
3 Nuclear masses in different macroscopic-microscopic models
In the last issue of the atomics mass table [53], there are 2259 measured and 906 estimated masses of isotopes with $Z,N\geq 20$ having an experimental error smaller than 1.5 MeV. All these masses are taken to find the best set of six adjustable parameters of our ISLD model (2). The mean square deviation of theoretical estimates and the experimental masses is minimized to obtain the best set of parameters. The microscopic energy $E_{\rm micr}$ from the Möller et al. mass-table [54] were used when evaluating the masses of isotopes using Eq. (4). Two parameter sets are found: one (a) corresponding to 2259 measured masses and the second one (b) obtained using 3165 measured and estimated masses denoted by the hash sign in the mass-table [53]. The r.m.s. deviation, which is a measure of the fit quality, is taken in the following form:
@@ 3 Nuclear masses in different macroscopic-microscopic models
Fig. 2: Difference of the atomic mass estimates obtained using the LSD [38] the ISLD(b) the LSD [38] models.
@@ 4 Potential energy surfaces and fission barrier heights
Fig. 4: Fission barrier heights evaluated using the topographical theorem [39] with the ISLD(a) (crosses), ISLD(b) (points), and LSD (open circles) set of parameters are compared with the empirical/experimental data taken from Ref. [58] (triangles) and [59] (stars) and with the theoretical values (squares) obtained within the mac-mic model in Ref. [60].
@@
The fission barrier heights estimated using Eq. (6) and the ISLD (a) (crosses) and (b) (circles) and the LSD (open circles) formulas are compared in Fig. 4 with the experimental data taken from Refs. [58] (triangles) and [59] (stars). In addition, the theoretical estimates of the highest fission barrier done in Ref. [60] are marked by squares. The experimental, or better to say empirical, fission barrier heights come mainly from analysis of the fission cross-section energy dependence and fissionability of nuclei. They are supplemented by data obtained from analysis of the excitation functions of spontaneously fissioning isomers and the group of strong resonances in the sub-barrier fission cross-section [58, 59]. The theoretical barrier heights tabulated in Ref. [60] were evaluated within the 7D mac-mic model with the YpE macroscopic energy part and the microscopic energy evaluated using the Woods-Saxon (WS) single-particle potential.
@@ 5 Conclusions
We have shown that the new liquid drop mass formula with the charge asymmetry term proportional to the isospin square (ISLD) describes the presently known experimental and estimated from systematic atomic masses well. One has to stress here that the ISLD formula contains only six adjustable parameters. The other models contain more free parameters, e.g., the twenty-one years old LSD mass formula [38], which reproduces the binding energies even better than the FRDM theory [54], possesses eight directly fitted parameters and four others (in the congruence/Wigner and odd-even terms) taken from the adjustments made in Ref. [9]. In this place, it is good to remember that microscopic models, typically heaving up to 14 interaction parameters, can reproduce 2149 experimental masses with the r.m.s. deviation equal to 0.798 MeV [61].
@@
It was also shown that both ISLD (b) and LSD models describe well the fission barrier heights of the heavy nuclei. The ISLD model predicts slightly larger atomic masses than the LSD one in the superheavy region of nuclei and for neutron-rich isotopes close to the neutron drip line. This difference in the mass estimates could be significant for SHN physics, and it can influence the nuclear r-process probabilities, which is very important in astrophysical theories. Also, the ISLD model predicts a more elongated shape of fissioning nuclei at scission configuration than the LSD formula. This effect can influence estimates of the fission fragment TKE and their charge equilibration (confer Ref. [62]).
We plan to perform extended dynamical calculations like those made in Refs. [49, 62, 63, 64, 65] but the use of the new ISLD formula when evaluating the potential energy surfaces of fissioning nuclei. The well-defined deformation dependence of the macroscopic energy term linear in $|N-Z|$ may also influence the isotopic yields of the fission fragments (confer Ref. [62]).

### v5

@@
Abstract   New liquid drop model with the iso-scalar volume and surface energy terms is applied to reproduce experimentally known masses of nuclei with number of protons and neutrons larger or equal to twenty. The ground-state microscopic energy corrections are taken into account. Even though the model contains only six adjustable parameters in the macroscopic part of the model, the quality of mass reproduction is good and comparable with other contemporary mass estimates. Also, the fission barrier heights of actinide nuclei evaluated using the topographical theorem of Myers and Świa̧tecki are close to the data.
Keywords: macro-micro model, nuclear masses, fission barrier heights, scission point properties
@@ 1 Introduction
Mass is the most fundamental property of atomic nuclei for its relevance in understanding not only the nuclear structure and the complex interaction in nuclei, but also the path of the synthesis of heavy nuclei in astrophysics [1]. In the super heavy region, the mass data also essential in the exploration of synthesizing super-heavy nuclide in terrestrial laboratory. Experimentally, the cooling store ring and the ion trap both provide precise measurement of nuclear mass in modern age [2, 3, 4, 5, 6]. So far, approximately 3000-4000 nuclides are measured while more than 9000 nuclides are believed to exist according to theoretic model calculations [7] . Descriptions and predictions of nuclear mass are of high theoretic interests and of significance in the calculation of reaction network in stellar environment including r-process [8]. Many models have been developed for this purpose. Roughly speaking, the models are classified in three groups. i) The macro-micro models, including finite range droplet model (FRDM) [9] and Weizsäcker-Skyrme (WS) models [10, 11, 12], ii) The pure micro models including Hartree-Fock-Bogoliubov (HFB) models [13], the covariant density-functional theory (CDFT) [14], relativistic Hartree-Bogoliubov theory [15, 16] etc. and iii) The local mass models which predict the atomic nucleus mass through the relationship of the masses of neighboring nuclides by taking into account different physical constraints [17, 18], such as Garvey-Kelson (GK) relations [19, 20] and the isobaric multiplet mass equation (IMME) [21] etc. Meanwhile, many efforts have been taken to evaluate and compare the predictive power of the models [22]. With the applications of the modern machine learning or neural network techniques, the predictive power of the models are further improved [23, 24, 25, 26, 27, 28]. Up to now, the residues of the nucleus mass reach the level of about 200 keV or better.
@@
A modern version of the LD model was proposed in 1966 by Myers and Świa̧tecki [34]. They have shown that the LD energy enriched by the shell and pairing effects can describe well the binding energies and quadrupole moments of known nuclei and gives a reasonable description of the fission-barrier height of heavy nuclei (see also [35]). Unfortunately, neither the Myers and Świa̧tecki LD formula nor its refined version called the droplet model could adequately reproduce the barrier heights of medium-heavy and lighter nuclei [36]. In addition, it was shown by von Groote and Hilf [37] that a further correction to the LD model, namely the curvature term, did not change much in this respect. Due to these results further development of the nuclear LD model was stopped in practice for more than three decades. Other much more complicated models like the droplet, Yukawa-folded (YF), Yukawa-plus-Exponential (YpE), or Finite-Range Droplet Models (FRDM) have been used to obtain within the so-called macro-micro approximation [34] of the binding energies and the fission barrier heights (for overview look, e.g., Ref. [38]). A more detailed historical review of the LD models can be found, e.g., in Ref. [39].
@@ 2 Theoretical model
$\begin{array}[]{ll}E_{\rm ISLD}(Z,A;{\rm sph})&=\displaystyle-a_{\rm vol}A\,\left(1-4\kappa_{\rm vol}\frac{T(T+1)}{A^{2}}\right)\\[12.91663pt] &\displaystyle+a_{\rm sur}A^{2/3}\,\left(1-4\kappa_{\rm sur}\frac{T(T+1)}{A^{2}}\right)\\[12.91663pt] &\displaystyle+{3\over 5}e^{2}\frac{Z(Z-1)}{A^{1/3}}+E_{\rm odd}(Z,A)\\[12.91663pt] &\displaystyle-{3\over 4}e^{2}\left({3\over 2\pi}\right)^{2/3}\frac{Z^{4/3}}{r_{0}A^{1/3}}~{},\end{array}$ (2)
@@
$E_{\rm odd}(Z,A)=\left\{\begin{array}[]{ll}{\cal D}/Z^{1/3}{}&{\rm for~{}Z~{}odd~{},}\\ {\cal D}/N^{1/3}{}&{\rm for~{}N~{}odd~{},}\\ {\cal D}/Z^{1/3}+{\cal D}/N^{1/3}{}&{\color[rgb]{0,0,0}\rm for~{}Z~{}and~{}N~{}odd~{},}\\ 0&{\rm for~{}Z~{}and~{}N~{}even~{}.}\end{array}\right.$ (3)
@@
Note that the linear in isospin term in the ISLD model (2) corresponds to the Wigner (or congruence) energy present in typical LD-like models (refer to, e.g., [34, 41]). In addition, in the ISLD model, the deformation dependence of this linear in $|N-Z|$ is well defined, whereas, in Ref. [41], one has to add an additional phenomenological deformation dependence in order to obtain a doubling of the congruence at the scission point, when two fission fragment nuclei are born. The only free, i.e., adjustable, parameters of the ISLD model are: $a_{\rm vol},\,\kappa_{\rm vol},\,a_{\rm sur},\,\kappa_{\rm sur},\,r_{0}$, and ${\cal D}$.
@@
and it is equal to the difference between the ground-state energy of the nucleus and the spherical macroscopic energy. The ground state microscopic energy corrections, take from the tables [56], were evaluated using the FRDM macroscopic deformation energy, the Yukawa-folded single particle potential, and the monopole pairing force.
Using the above microscopic energy correction from the tables [56], we make an approximation. Namely, we assume that the stiffness of the macroscopic deformation energy evaluated within the FRDM is close the that obtained using the ISLD model. A detailed calculation shows that the above approximation may lead to around 50 keV inaccuracy of our mass estimates, which is about ten times smaller than the r.m.s. deviation of the fitted masses shown in Tab. 2. In adition, a similar way of evaluating the nuclear binding energies was also used in Refs. [41, 40, 46].
The Fourier-over-Spheroid (FoS) shape parametrization [50, 51] was used to describe the shape of fissioning nuclei
$\rho_{s}^{2}(z)=\frac{R_{0}^{2}}{c}\,f\left(\frac{z-z_{\rm sh}}{z_{0}}\right)~{}.$ (6)
where the function $f(u)$ with $u\in[-1,1]$ is defined as follows:
$f(u)=1-u^{2}-\sum\limits_{k=1}^{n}\left\{a_{2k}\cos({2k-1\over 2}\pi u)+a_{2k+1}\sin(k\pi u)\right\}$ (7)
Here $\rho_{s}(z)$ is the distance of a surface point to the $z$-axis and $z_{0}=cR_{0}$, with $R_{0}$ being the sphere’s radius, is the half-length of the deformed nucleus. The first two terms in $f(u)$ describe a spheroid, while the others give the deviation of the nucleus surface from the spheroidal form. The shift parameter $z_{\rm sh}=-3/(4\pi)\,z_{0}(a_{3}-a_{5}/2+\dots)$ ensures that the origin of the coordinate system is located at the center of mass of the nucleus. The volume conservation condition implies $a_{2}=a_{4}/3-a_{6}/5+\dots$. The expansion coefficients $a_{i}$ are treated as the deformation parameters. The parameter $c$ determines the elongation of the nucleus, keeping its volume fixed, while $a_{3}$ and $a_{4}$ are, respectively, the deformation parameters essentially responsible for the reflection asymmetry and the neck formation of the deformed shape.
The FoS parametrization reproduces nuclear shapes very close to the optimal shapes obtained by the Strutinsky variational procedure [52] and allows to evaluate precisely the liquid drop energy at the saddle point. Knowing this energy, one can estimate the fission barrier height ($V_{\rm sadd}$) with the help of the Myers and Świa̧tecki topographical theorem [41]
@@
3 Nuclear masses in different macro-micro models
In the last issue of the atomics mass table [55], there are 2259 measured and 906 estimated masses of isotopes with $Z,N\geq 20$ having an experimental error smaller than 1.5 MeV. All these masses are taken to find the best set of six adjustable parameters in the marcroscopic part of our ISLD model (2). The mean square deviation of theoretical estimates and the experimental masses is minimized to obtain the best set of parameters. The microscopic energy $E_{\rm micr}$ from the Möller et al. mass-table [56] were used when evaluating the masses of isotopes using Eq. (4). Two parameter sets are found: one (a) corresponding to 2259 measured masses and the second one (b) obtained using 3165 measured and estimated masses denoted by the hash sign in the mass-table [55]. The r.m.s. deviation, which is a measure of the fit quality, is taken in the following form:
@@ 3 Nuclear masses in different macro-micro models
Fig. 2: Difference of the atomic mass estimates obtained using the ISLD(b) and the LSD [40] models.
@@ 4 Potential energy surfaces and fission barrier heights
Fig. 4: Fission barrier heights evaluated using the topographical theorem (see Eq. 8) with the ISLD(a) (crosses), ISLD(b) (points), and LSD (open circles) set of parameters are compared with the empirical/experimental data taken from Ref. [60] (triangles) and [61] (stars) and with the theoretical values (squares) obtained within the macro-micro model in Ref. [62].
@@
The fission barrier heights estimated using Eq. (8) and the ISLD (a) (crosses) and (b) (circles) and the LSD (open circles) formulas are compared in Fig. 4 with the experimental data taken from Refs. [60] (triangles) and [61] (stars). In addition, the theoretical estimates of the highest fission barrier done in Ref. [62] are marked by squares. The experimental, or better to say empirical, fission barrier heights come mainly from analysis of the fission cross-section energy dependence and fissionability of nuclei. They are supplemented by data obtained from analysis of the excitation functions of spontaneously fissioning isomers and the group of strong resonances in the sub-barrier fission cross-section [60, 61]. The theoretical barrier heights tabulated in Ref. [62] were evaluated within the 7D macro-micro model with the YpE macroscopic energy part and the microscopic energy evaluated using the Woods-Saxon (WS) single-particle potential.
@@ 5 Conclusions
We have shown that the new liquid drop mass formula with the charge asymmetry term proportional to the isospin square (ISLD) describes the presently known experimental and estimated from systematic atomic masses well. One mast stress that the ISLD formula for the macroscopic energy contains only six adjustable parameters. The other models contain more free parameters, e.g., the twenty-one years old LSD mass formula [40], which reproduces the binding energies even better than the FRDM theory [56], possesses eight directly fitted parameters and four others (in the congruence/Wigner and odd-even terms) taken from the adjustments made in Ref. [9]. In this place, it is good to remember that microscopic models, typically having up to 14 interaction parameters, can reproduce 2149 experimental masses with the r.m.s. deviation equal to 0.798 MeV [63].
@@
It was also shown that both ISLD (b) and LSD models describe well the fission barrier heights of the heavy nuclei. The ISLD model predicts slightly larger atomic masses than the LSD one in the superheavy region of nuclei and for neutron-rich isotopes close to the neutron drip line. This difference in the mass estimates could be significant for SHN physics, and it can influence the nuclear r-process probabilities, which is very important in astrophysical theories. Also, the ISLD model predicts a more elongated shape of fissioning nuclei at scission configuration than the LSD formula. This effect can influence estimates of the fission fragment TKE and their charge equilibration (refer to Ref. [64]).
We plan to perform extended dynamical calculations like those made in Refs. [51, 64, 65, 66, 67] but the use of the new ISLD formula when evaluating the potential energy surfaces of fissioning nuclei. The well-defined deformation dependence of the macroscopic energy term linear in $|N-Z|$ may also influence the isotopic yields of the fission fragments (refer to Ref. [64]).